Check whether the sequence, {Sn} where

Sn=1/1!+1/3!+1/5!+....1/(2n+1)!
Is convergent or not.
1
Expert's answer
2020-02-27T09:44:35-0500

"S=\frac{1}{1!}+\frac{1}{3!}+\frac{1}{5!}+...+\frac{1}{(2n+1)!}+..."

apply the d'Alembert ratio test to

"a_n=\frac{1}{(2n+1)!},\\
a_{n+1}=\frac{1}{(2n+3)!},\\"

consider

"lim_{n\mapsto \infty} \frac{a_{n+1}}{a_n}=lim_{n\mapsto \infty} \frac{(2n+1)!}{(2n+3)!}=\\
=lim_{n\mapsto \infty} \frac{(2n+1)!}{(2n+1)!(2n+2)(2n+3)}=\\
=lim_{n\mapsto \infty} \frac{1}{(2n+2)(2n+3)}=0<1."

Since the limit value 0 is less than 1, the series is convergent.